Could you simply 16 1/2 and describe how you got your answer?
Answer by stanbon(75887)

simply 16 1/2 and describe how you got your answer?
-------
Ans: (32+1)/2 = 33/2
--------
Procedure: the number of (1/2) in 16 is 2*16 = 32
---
Add 32/2 to 1/2 to get 33/2
==============================
Cheers,
Stan H.
=============